OpenStudy (anonymous):

An experiment consists of selecting a number at random between 0 and 9 inclusive, and a letter from the English alphabet. What is the size of the sample space of this experiment? 35 36 52 234 260

OpenStudy (anonymous):

WILL GIVE MEDAL :) PLEASE HELP

OpenStudy (anonymous):

@iGreen @shinalcantara @felavin

OpenStudy (anonymous):

The sample space is {0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 0, a, b, c, d, e, f, g, ... , z} So the size of the sample space is 10 + 26 = 36 Answer : B

OpenStudy (anonymous):

\(X = \{0, ..., 9\}\) and \(Y = \{a, .., z\}\) The sample space is then \(S = X \times Y\) Therefore we have that \(|S| = |X||Y| = 10 \cdot 26 = 260\)

OpenStudy (anonymous):

A sample for the experiment includes both a number AND letter.

OpenStudy (anonymous):

Each trial of the experiment gives you both a number and letter. So a sample is an ordered pair \((x, y) \in \{0, ..., 9\} \times \{a, ..., z\}\)